# Exploration and Incentivizing Participation in Clinical Trials

### v8

@@ Abstract
Participation incentives a well-known issue inhibiting randomized clinical trials (RCTs). We frame this issue as a non-standard exploration-exploitation tradeoff: an RCT would like to explore as uniformly as possible, whereas each patient prefers “exploitation”, i.e., treatments that seem best. We incentivize participation by leveraging information asymmetry between the trial and the patients. We measure statistical performance via worst-case estimation error under adversarially generated outcomes, a standard objective for RCTs. We obtain a near-optimal solution in terms of this objective: an incentive-compatible mechanism with a particular guarantee, and a nearly matching impossibility result for any incentive-compatible mechanism. We consider three model variants: homogeneous patients (of the same “type” comprising preferences and medical histories), heterogeneous agents, and an extension with estimated type frequencies.
@@ 1 Introduction
Randomized Clinical Trials (RCTs) are a standard way to evaluate safety and efficiency of medical treatments. [^1] We use “RCT” acronym in this sense throughout. Another common meaning, “randomized controlled trial”, is also consistent with our model, since one of the arms there is typically a “control”. A paradigmatic design considers $n=2$ alternative treatments, a.k.a. arms, e.g., a new and an old drug for the same medical condition, or a new drug and a placebo. Each patient $t$ is assigned an arm $a_{t}$ independently and uniformly at random. An outcome $\omega_{t}$ is observed and evaluated according to some predefined metric $f(\cdot)$. The goal is to use the observations to estimate the counterfactual averages $\tfrac{1}{T}\,\sum_{t\in[T]}f(\omega_{a,t})$, $a\in[n]$, where $T$ is the number of patients. Independent per-patient randomization over arms is essential to remove selection bias and ensure statistical validity. Uniform randomization is usually preferred, as it minimizes the estimation error. In particular, one can obtain strong provable guarantees even if each arm’s outcomes are generated by an adversary, rather than sampled from some fixed (arm-dependent) distribution. [^2] This corresponds to (oblivious) adversary in adversarial bandits (Auer et al.,, 2002).
@@
A well-known issue which inhibits RCTs is the difficulty of recruiting patients (e.g., see surveys Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). This issue may be particularly damaging for large-scale RCTs for wide-spread medical conditions with relatively inexpensive treatments. Then finding suitable patients and providing them with appropriate treatments would be fairly realistic, but incentivizing patients to participate in sufficient numbers may be challenging.
We focus on patients’ reluctance to participate that stems from their preferences among the available treatments (as opposed to other factors, e.g., logistical difficulties). We frame this issue as a non-standard variant of exploration-exploitation tradeoff: the RCT always prefers to explore as uniformly as possible, whereas each patient prefers to exploit (choose treatment that seems best), and might not participate otherwise.  [^3] The standard variant of this tradeoff has an algorithm balance exploration vs exploitation for its own sake, so as to converge on the best arm.
@@
We initiate the study of incentivized participation in RCTs. Our objective is to incentivize participation while optimizing statistical performance. We leverage information asymmetry between the RCT and the patients. We use preliminary data from the ongoing trial to skew the random choice of arms in favor of a better treatment (i.e., towards exploitation), while retaining a sufficient amount of near-uniform exploration. This causes the patients to prefer participation to any alternative treatment options they might have (which also don’t have access to the trial’s data).
@@
We make two assumptions that are common in economic theory: the RCT has the power to (i) prevent undesirable information disclosure to the patients, and (ii) commit to following a particular policy. While such assumptions can be problematic in other applications, they are realistic in RCTs. Indeed, it is a standard practice to not reveal the history of an ongoing RCT and the assigned treatment, to avoid biasing patients and doctors in favor of better-performing treatments (e.g., see Detry et al., (2012, pp. 26-30)). Moreover, the design of an RCT is typically required by regulation to be fixed in advance, spelling out all applicable “decision points”, and (essentially) it needs to be revealed to the patients to meet the standards of “informed consent” (Arango et al.,, 2012).
@@
While RCTs with $n=2$ arms is the paradigmatic case, our results seamlessly extend to $n>2$. RCTs with $\geq 3$ arms are studied in biostatistics, e.g., (Hellmich,, 2001; Freidlin et al.,, 2008), and are important in practice, e.g., (Parmar et al.,, 2014; Redig and Jänne,, 2015).
@@
Our modelling approach. We model “incentivized participation” as a mechanism design problem under a statistical objective and incentives constraints. An RCT (mechanism) sequentially interacts with self-interested patients (agents) like a multi-armed bandit algorithm: in each round, an agent arrives, the mechanism chooses an arm for this agent, and observes an outcome. The mechanism’s statistical objective is standard for RCTs: counterfactually estimate each arm’s performance on the whole population, even if the outcomes are generated by an adversary. [^4] In contrast, the standard objective for multi-armed bandits is to optimize cumulative payoffs. If not for the incentives constraints, the mechanism would simply choose arms uniformly at random.
@@
Crucially, the agents believe in much simpler world than the adversarial one allowed by our statistical model. Essentially, they believe one’s outcome distribution is determined by the assigned arm and the medical history. Such beliefs are fairly realistic, representing pre-existing medical knowledge, and appear necessary for tractability.
@@
We consider three model variants. To isolate the core issues of incentivized participation, we consider on the fundamental case of homogeneous agents with the same preferences and medical histories. The general case of heterogeneous agents endows each agent with a type: a (public) medical history which determines the outcome distributions and (private) preferences over these outcomes. The third variant provides the mechanism with estimated type frequencies; this allows to mitigate the harmful influence of rare-but-difficult agent types.
@@
Our results. We focus on mechanisms with simple two-stage designs: a relatively short warm-up stage of predetermined duration, followed by the main stage that needs to be counterfactually estimated. The data from the warm-up stage creates information asymmetry which drives participation incentives, whereas the data from the main stage determines the counterfactual estimates. [^5] In particular, the warm-up data is discarded for the counterfactual estimates. This is a part the price of incentivizing participation in a (much larger) main stage.
@@
Our guarantees focus on the main stage, assuming that enough data is collected in the warm-up stage. The aforementioned data can also be collected either endogenously (i.e., by a mechanism in our model), or exogenously, e.g., using paid volunteers. The key point is that a relatively small amount of data, determined by the prior and the utility structure, suffices to bootstrap the main stage for an arbitrarily large number of agents. Endogenous implementation of the warm-up stage can rely on prior work on incentivized exploration (more on this in “Related work” and Lemma 3.6).
@@
We obtain a near-optimal solution for each model variants described above, from homogeneous agents to heterogeneous agents to estimated type frequencies. First, we put forward a benchmark: a hypothetical policy that observes the true “state of the world” according to agents’ beliefs [^6] A mapping from arms (or type-arm pairs) to outcome distributions. and optimizes a certain objective. Then, we design a mechanism that attains a worst-case upper bound on the estimation error, stated in terms of this benchmark and scaling as $1/T$. Finally, we prove that no mechanism achieves better estimation error, up to constant factors. Conceptually, we connect the dynamic process and statistical estimators with a static information-design problem. Many ideas carry over from one variant to the next.
@@
Our mechanisms are simple in that their “sampling distributions” (the distributions from which the arms are sampled) do not adapt to the data collected in the main stage. This is desirable in practice.777 Because treating the patients and/or observing the outcomes may take a long time, and the doctors involved may find it difficult to coordinate with one another. Consequently, many/most patients may need to be treated in parallel. So, RCTs typically do not adapt to data, or have just a few pre-determined “review points” when one can make a data-adaptive decision, with non-data-adaptive “stages” in between. In contrast, our negative results apply to a wide class of mechanisms, including the “multi-stage” mechanisms from Footnote 7. In particular, the sampling distributions can change over time and in “almost all” rounds they can depend arbitrarily on all past observations.888“Almost all” rounds here means: all rounds outside of a contiguous time interval comprising a constant (but possibly very small) fraction of rounds.
@@
Discussion. Our contribution is not the mechanisms per se, but the whole “journey”: from the new model to the benchmark to the matching upper/lower bounds, and from homogeneous agents to heterogeneous agents to estimated type frequencies. While our results are asymptotic, and hence most meaningful for RCTs with a large number of participants, participation incentives would make such trials more common and realistic. For concreteness, we also illustrate our results with simple numerical examples (Appendix A).
@@
We emphasize that theoretical participation incentives come at a price: the warm-up samples are discarded for counterfactual estimates, and the main-stage sampling distributions are skewed. Whether this price is acceptable in practice depends on how much these incentives actually help in recruiting patients. Ultimately, the “price of incentives” should be included in the “price of experimentation” that the society pays for the sake of obtaining reliable medical knowledge.
@@
This paper is only concerned with patients’ initial decision to participate in an RCT. Their incentives to remain in the trial, e.g., if the treatment appears ineffective, is outside of our scope. [^9] This is indeed an issue in practice, as per (FDA,, 2001, Section 2.1.7.2). However, it is unclear if any approach based on information asymmetry could possibly help here: once the patient learns about the actual outcome, the “principal” does not have much/any informational advantage left.
Map of the paper. Section 2 spells out our model, focusing on the homogeneous agents. Section 3 works out the case of homogeneous agents, introducing key ideas with a more accessible exposition. Heterogeneous agents are treated in Section 4; Section 5 considers an extension with estimated type frequencies. The technical machinery in Sections 3, 4 and 5 is introduced gradually, and heavily reused in the later sections.
@@ 1.1 Related work
Incentivized exploration (IE) is a closely related model, motivated by recommendation systems. The platform strives to incentivize users to explore when they would prefer to exploit. The platform issues recommendations (which the users can override), and makes them incentive-compatible via information asymmetry. Introduced in (Kremer et al.,, 2014; Che and Hörner,, 2018), IE is well-studied, see (Slivkins,, 2019, Chapter 11) for a survey. Most technically related are Mansour et al., (2020); Immorlica et al., (2019).
The main difference concerns performance objectives. IE strives to maximize the total reward under time-invariant outcome distribution, or just explore each arm once. Neither serves our statistical objective, which requires independent near-uniform randomization in each round: reward-maximization pulls a mechanism in a very different direction, and merely exploring each arm $N$ times does not ensure statistical validity under adversarially chosen outcomes. The economic constraints are different, too, because patients cannot observe or alter their treatments (but can choose whether to participate). [^10] This distinction is consequential only for $n>2$ arms. Then the economic constraint in IE is stronger than ours. Mechanism’s power to control information and commit to a particular policy are substantial assumptions in IE (e.g., Bahar et al., (2016); Immorlica et al., (2020) are written to mitigate them), but they are very benign for RCTs, as discussed above. Hence, our mechanisms, analyses and technical guarantees are very different from those in IE. However, our mechanisms reuse IE for the “warm-up stage” (see Lemma 3.6).
@@
Randomized Clinical Trials (RCTs). We only point out the key connections to the (huge) medical literature on RCTs, as a more detailed review is beyond our scope.
@@
2 Our model: incentivized clinical trial
We first present our model for homogeneous agents (i.e., agents with the same “type”), then extend it to heterogeneous agents in Section 4. This is to disentangle the core model, which is quite complex even for a single agent type, from some subtlety and heavy notation needed for multiple types.
Thus, we consider a randomized clinical trial (RCT) with some patients, interpreted as a mechanism) and agents, respectively. The mechanism interacts with the agents as a multi-armed bandit algorithm, but optimizes a very different statistical objective, and operates under economic constraints to incentivize participation. Accordingly, we split the model description into three subsections: the bandit model for the interaction protocol the statistical model for its performance objective, and the economic model for the agents’ incentives.
@@ 2.1 Bandit model: the interaction protocol
Remark 2.1.
The agents cannot change the assigned arms. Indeed, Indeed, an RCT participant cannot change her assigned treatment and cannot even observe what this treatment is.
@@ 2.2 Statistical model: the performance objective
The mechanism should estimate the average score of each arm without any assumptions on the adversary, i.e., on how the outcomes are generated. Not relying on modeling assumptions is a standard desiderata in randomized clinical trials.
@@ 2.3 Economic model: agents, beliefs and incentives
The agents’ information structure is as follows. (i) Each agent $t$ does not observe anything about the previous rounds. Indeed, not revealing any information about an ongoing trial is strongly encouraged in practice (e.g.,  Arango et al.,, 2012). Consequently, agent $t$ does not observe the chosen distribution $p_{t}$ if it depends on the history. (ii) Each agent $t$ does not observe the assigned arm $a_{t}$. Indeed, RCTs typically present different treatments in an identical way, e.g., as a pill that looks and feels exactly the same; this is strongly encouraged whenever possible (e.g.,  Arango et al.,, 2012). (iii) Each agent knows the specification of the mechanism before (s)he arrives. In practical terms, the entity which runs the RCT (e.g., a hospital) commits to a particular design of the trial, and this design is revealed to all patients. This property is typically enforced via regulation.
The mechanism must incentivize participation, i.e., satisfy Individual Rationality (IR). We require Bayesian IR in expectation over the agents’ beliefs over the adversaries. [^14] Satisfying the IR property for every adversary (ex-post IR) is clearly impossible, except for a trivial mechanism that always chooses the agent’s outside option. Crucially, we posit that agents’ beliefs are over stochastic adversaries, i.e., the agents believe that a given treatment yields the same outcome distribution in all patients. [^15] Agents have misspecified beliefs, in the sense that their support does not necessarily include the actual adversary. In particular, the actual adversary need not be stochastic. Misspecified beliefs are studied in a considerable recent literature in economic theory (starting from Berk,, 1966; Esponda and Pouzo,, 2016). However, this literature tends to focus on convergence and equilibria (e.g., Fudenberg et al.,, 2021; Esponda and Pouzo,, 2021; Esponda et al.,, 2021), whereas our paper takes the mechanism design perspective and is concerned with IPS-style statistical estimation. All agents share the same belief, a common simplification in economic theory. Informally, the beliefs represents the current medical knowledge.
Let us specify the agents’ beliefs. The outcome of each arm $a$ is an independent draw from some distribution $\psi^{*}(a)$ over outcomes each time this arm is chosen. For better notation, we identify states as mappings $\psi:[n]\to\Delta_{\Omega}$ from arms to distributions over outcomes, and $\psi^{*}$ as the true state. [^16] We use a standard notation: $\Delta_{\Omega}$ is the set of all distributions over $\Omega$. Then, the agents believe that $\psi^{*}$ is drawn from some Bayesian prior: distribution $\mathcal{P}$ over states. We posit that $\mathcal{P}$ is known to the mechanism, and assume that it has finite support.
@@
We ensure that each agent weakly prefers participation to the “outside option”. To formalize the latter, we posit that an agent could refuse to participate in the mechanism and choose an arm that appears best based on the initial information. This achieves Bayesian-expected utility
$\textstyle U_{\mathtt{out}}:=\max_{a\in[n]}\;\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\;\left[\,U_{\psi}(a)\,\right].$ (2.3)
@@
While (2.3) is a natural “outside option” within our model, our analysis accommodates an arbitrary, exogenously given $U_{\mathtt{out}}\in[0,1]$, and all results carry over with minimal modifications. This allows us to model several effects. First, some medical treatments might not be available outside of a clinical trial. Then we could take the $\max$ in (2.3) only over the treatments that are. Conversely, a standard treatment might not be available inside the trial, e.g., in a new treatment vs. placebo trials. Then the $\max$ in (2.3) should also include the standard treatment. Finally, patients might assign positive utility to participation itself, e.g., because they value contributing to science or they expect to receive higher-quality and/or less expensive medical care. This can be modeled (albeit in a rather idealized way) by decreasing $U_{\mathtt{out}}$ in (2.3) by some known amount.

### v9

# Exploration and Incentivized Participation in Randomized Trials

@@ Abstract
Participation incentives is a well-known issue inhibiting randomized controlled trials (RCTs) in medicine, as well as a potential cause of user dissatisfaction for RCTs in online platforms. We frame this issue as a non-standard exploration-exploitation tradeoff: an RCT would like to explore as uniformly as possible, whereas each “agent” (a patient or a user) prefers “exploitation”, i.e., treatments that seem best. We incentivize participation by leveraging information asymmetry between the trial and the agents. We measure statistical performance via worst-case estimation error under adversarially generated outcomes, a standard objective for RCTs. We obtain a near-optimal solution in terms of this objective: an incentive-compatible mechanism with a particular guarantee, and a nearly matching impossibility result for any incentive-compatible mechanism. We consider three model variants: homogeneous agents (of the same “type” comprising beliefs and preferences), heterogeneous agents, and an extension with estimated type frequencies.
Keywords: exploration-exploitation tradeoff, incentive-compatiblity, economic design, randomized controlled trials, clinical trials, multi-armed bandits
@@ 1 Introduction
Randomized Controlled Trials (RCTs) is a gold-standard way to evaluate safety and efficiency of medical treatments. A paradigmatic design considers $n=2$ alternatives, a.k.a. arms. [^2] While one arm is typically a “control group” (such as an old treatment or a placebo), this distinction is not essential throughout this paper. Each patient $t$ is assigned an arm $a_{t}$ independently and uniformly at random. An outcome $\omega_{t}$ is observed and evaluated according to some predefined metric $f(\cdot)$. The goal is to use the observations to estimate the counterfactual averages $\tfrac{1}{T}\,\sum_{t\in[T]}f(\omega_{a,t})$, $a\in[n]$, where $T$ is the number of patient. Independent per-patient randomization over arms is essential to remove selection bias and ensure statistical validity. Uniform randomization is usually preferred, as it minimizes the estimation error. In particular, one can obtain strong provable guarantees even if each arm’s outcomes are generated by an adversary, rather than sampled from some fixed (arm-dependent) distribution. [^3] In the same sense as “oblivious adversary” in adversarial bandits (Auer et al.,, 2002).
@@
A well-known issue which inhibits RCTs is the difficulty of recruiting patients (e.g., see surveys Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). This issue may be especially damaging for large-scale RCTs for wide-spread medical conditions with relatively inexpensive treatments. Then finding suitable patients and providing them with appropriate treatments would be fairly realistic, but incentivizing patients to participate in sufficient numbers may be challenging.
Similar issues arise in user-facing online platforms, where experimentation is pervasive and RCTs, a.k.a. A/B testing, is a standard approach. [^4] For background on A/B testing, see e.g., Kohavi et al., (2009); Kohavi and Longbotham, (2017). While platform users are typically not informed of the particular experiments they are assigned to, the loss in performance (whether actual or perceived) may lead to user dissatisfaction. Having robust participation incentives would therefore be reassuring to the users, and may allow for more ambitious and risky experiments.
The standard approach of subsidizing the agents may be too expensive at the scale required. Indeed, clinical trials may require large subsidies when health outcomes are at stake, and online platforms may need a very large number of participants. The subsidies may also lead to selection bias towards poorer or more adventurous participants.
Our scope. We initiate the study of incentivized participation in RCTs. Our objective is to incentivize participation while optimizing statistical performance. We frame it as a non-standard variant of exploration-exploitation tradeoff: the RCT always prefers to explore as uniformly as possible, whereas each patient/user prefers to exploit (choose an arm that seems best), and might not participate if the RCT is inconsistent with this preference. [^5] The standard tradeoff has an algorithm balance exploration vs exploitation to converge on the best arm. We henceforth refer to the potential RCT participants as agents.
@@
We leverage information asymmetry between the RCT and the agents. We use preliminary data from the ongoing trial to skew the random choice of arms in favor of a better treatment (i.e., towards exploitation), while retaining a sufficient amount of near-uniform exploration. This causes the agents to prefer participation in the RCT to the alternatives available externally.
@@
We make two standard assumptions from economic theory: the RCT has the power to (i) prevent undesirable information disclosure to the agents, and (ii) commit to following a particular policy. These assumptions are realistic in clinical RCTs due to regulation. Indeed, the history of an ongoing RCT and the assigned treatment should not revealed to the patients (or their doctors) to the extent possible, to avoid biasing them in favor of better-performing treatments (Detry et al.,, 2012, pp. 26-30). [^6] In particular, different treatments are typically presented in the same way, e.g., as externally-similar pills. Moreover, the design of an RCT must be fixed in advance, spelling out all applicable decision points, and revealed to the patients (Arango et al.,, 2012). Experimentation in online platforms is typically not revealed to the users, and the commitment power is a standard assumption in the literature (e.g., see “incentivized exploration” in Related Work).
@@
While RCTs with $n=2$ arms is the paradigmatic case, our results seamlessly extend to $n>2$. Note that clinical RCTs with $\geq 3$ arms are studied in biostatistics, e.g., (Hellmich,, 2001; Freidlin et al.,, 2008), and are important in practice, e.g., (Parmar et al.,, 2014; Redig and Jänne,, 2015).
@@
Our modelling approach. We model “incentivized participation” as a mechanism design problem under a statistical objective and incentives constraints. An RCT (mechanism) sequentially interacts with self-interested agents like a multi-armed bandit algorithm: in each round, an agent arrives, the mechanism chooses an arm for this agent, and observes an outcome. The mechanism’s statistical objective is standard for RCTs: counterfactually estimate each arm’s performance on the whole population, even if the outcomes are generated by an adversary. [^7] In contrast, the standard objective for multi-armed bandits is to optimize cumulative payoffs. Formally, the mechanism optimizes the worst-case estimation error for these counterfactual estimates, denoted $\mathtt{ERR}$. If not for the incentives constraints, the mechanism would simply choose arms uniformly at random.
Agents have utilities over the outcomes, possibly different from the “societal value” estimated by the RCT. The RCT may also be interested in the empirical frequency of a particular outcome (we allow a generic formulation which captures both societal value and outcome frequency).
@@
While our statistical guarantees must be robust against adversarial outcomes, the outcome-generating process could be much more well-behaved. Crucially, the agents believe that it is. Specifically, the agents believe that one’s outcome distribution is determined by the assigned arm and one’s observable attributes (resp., medical history or user’s profile). While such beliefs may be incorrect, they appear necessary for tractability, and are particularly appropriate in clinical RCTs as a way to represent pre-existing medical knowledge.
@@
We consider three model variants. To isolate the core issues of incentivized participation, we consider on the fundamental case of homogeneous agents with the same beliefs and preferences. The general case of heterogeneous agents endows each agent with a type: observable attributes which determine the outcome distributions in the agent’s beliefs and unobservable (private) preferences over the outcomes which need to be elicited in an incentive-compatible way. [^8] In practice, preference elicitation would happen during registration/opt-in process. For online platforms, private preferences and their elicitation make sense only if the RCT has explicit opt-in. The latter is uncommon, but may be needed if the RCT involves high-stakes interactions (such as buying a house) or is potentially controversial. The third variant provides the mechanism with estimated type frequencies over the agent population; this allows to mitigate the harmful influence of rare-but-difficult agent types.
@@
Our results. Our mechanisms have simple two-stage designs: a relatively short warm-up stage of predetermined duration, followed by the main stage that needs to be counterfactually estimated. The data from the warm-up stage (warm-up data) creates information asymmetry in the main state which drives participation incentives, whereas the main stage determines the counterfactual estimates. A warm-up stage which influences the main stage is very common in practice. [^9] E.g., phase-I clinical trials to calibrate the dosage and “pilot experiments” in online platforms.
@@
Our guarantees focus on the main stage, given enough warm-up data. A relatively small amount of warm-up data, determined by the prior and the utility structure, suffices to bootstrap the main stage for an arbitrarily large number of agents. The warm-up data can be collected exogenously e.g., using paid volunteers, or endogenously by a mechanism in our model, e.g., via prior work on incentivized exploration (see “Related work” and Lemma 3.6).
@@
We obtain a near-optimal solution for each model variants described above, from homogeneous agents to heterogeneous agents to estimated type frequencies. First, we put forward a benchmark: a hypothetical policy that observes the true “state of the world” according to agents’ beliefs and optimizes a certain objective. Then, we design a mechanism whose estimation error ($\mathtt{ERR}$) is upper-bounded in terms of this benchmark. Finally, we prove that no mechanism achieves better estimation error, up to constant factors. Many ideas carry over from one model variant to the next.
@@
Our mechanisms have a desirable property that their “sampling distributions” (the distributions from which the arms are sampled) do not adapt to the data collected in the main stage. Indeed, clinical RCTs in practice typically do not adapt to data, except possibly at a few pre-determined “review points”,101010 This is because treating the patients and/or observing the outcomes may take a long time, and the doctors may find it difficult to coordinate across the patients. So, many/most patients may need to be treated in parallel. and neither do most RCTs in online platforms due to engineering considerations. In contrast, our negative results apply to a wide class of mechanisms, including the “few review points” mechanisms mentioned above.
Our contribution is not the mechanisms per se, but the whole “journey”: from the new model to the benchmark to the matching upper/lower bounds, and from homogeneous agents to heterogeneous agents to estimated type frequencies. Conceptually, we connect the dynamic process and statistical estimators in the mechanism with a static information-design problem in the benchmark.
@@
Additional discussion on clinical RCTs. While our results are asymptotic, and hence most meaningful for RCTs with a large number of patients, participation incentives would make such trials more common and realistic. For concreteness, we provide numerical examples (Appendix A).
@@
Theoretical participation incentives come at a price: the warm-up samples are discarded for counterfactual estimates, and the main-stage sampling distributions are skewed. This “price of incentives” should be included in the overall “price of experimentation” incurred to obtain reliable medical knowledge. Whether it is acceptable in practice depends on how many patients these incentives actually help to recruit.
@@
We focus on patients’ reluctance to participate that stems from their preferences among the available treatments (as opposed to other factors, e.g., logistical difficulties). Further, this paper is only concerned with patients’ initial decision to participate in an RCT. Their incentives to remain in the trial, e.g., if the treatment appears ineffective, is outside of our scope. [^11] This is indeed an issue in practice, as per (FDA,, 2001, Section 2.1.7.2). However, it is unclear if any approach based on information asymmetry could possibly help here: once the patient learns about the actual outcome, the “principal” does not have much/any informational advantage left.
@@ 1.1 Related work
Incentivized exploration (IE) is a closely related model concerning experimentation on online platforms. Here the “platform” is a recommendation system which strives to incentivize users to explore when they would prefer to exploit. The platform issues recommendations (which the users can override), and makes them incentive-compatible via information asymmetry. Introduced in (Kremer et al.,, 2014; Che and Hörner,, 2018), IE is well-studied, see (Slivkins,, 2019, Chapter 11) for a survey. Most technically related are Mansour et al., (2020); Immorlica et al., (2019).
The main difference concerns performance objectives. IE strives to maximize the total reward under time-invariant outcome distribution, or just explore each arm once. Neither serves our statistical objective, which requires independent near-uniform randomization in each round: reward-maximization pulls a mechanism in a very different direction, and merely exploring each arm $N$ times does not ensure statistical validity under adversarially chosen outcomes. The economic constraints are different, too, because agents in our model can only choose whether to participate. [^12] This distinction is consequential only for $n>2$ arms. Then the economic constraint in IE is stronger than ours. Hence, our mechanisms, analyses and technical guarantees are very different from those in IE. However, our mechanisms reuse IE for the “warm-up stage” (see Lemma 3.6).
@@
Clinical RCTs. We only point out the key connections to the (huge) medical literature on RCTs, as a more detailed review is beyond our scope.
@@
1.2 Map of the paper
We first present and analyze our model for the case of homogeneous agents (Sections 2 and 3). This is to introduce the core model and key ideas in a more accessible way, disentangling them from the additional subtlety and heavier notation needed for the general case. Heterogeneous agents are defined and treated in Section 4. Section 5 works out an extension with estimated type frequencies. The technical machinery in Sections 3, 4 and 5 is introduced gradually and heavily reused.
2 Our model: incentivized RCT (homogeneous agents)
We focus on homogeneous agents, i.e., agents with the same “type”, as stated in Section 1.2. We consider a randomized controlled trial (RCT) with some patients/users, respectively interpreted as a mechanism and agents. The mechanism interacts with the agents as a multi-armed bandit algorithm, but optimizes a very different statistical objective, and operates under economic constraints to incentivize participation. Accordingly, we split the model description into three subsections: the bandit model for the interaction protocol the statistical model for its performance objective, and the economic model for the agents’ incentives.
@@ 2.2 Statistical model: the performance objective
The mechanism should estimate the average score of each arm without any assumptions on the adversary, i.e., on how the outcomes are generated. Not relying on modeling assumptions is a standard desiderata in randomized controlled trials.
@@ 2.3 Economic model: agents, beliefs and incentives
The agents’ information structure is as follows (as motivated in the Introduction). (i) Each agent $t$ does not observe anything about the previous rounds. Consequently, agent $t$ does not observe the chosen distribution $p_{t}$ if it depends on the history. (ii) Each agent $t$ does not observe the assigned arm $a_{t}$. (iii) Each agent knows the specification of the mechanism before (s)he arrives.
The mechanism must incentivize participation, i.e., satisfy Individual Rationality (IR). We require Bayesian IR in expectation over the agents’ beliefs over the adversaries. [^17] We invoke agents’ beliefs because satisfying the IR property for every adversary (a.k.a. ex-post IR) is clearly impossible, except for a trivial mechanism that always chooses the agent’s outside option. Crucially, we posit that agents’ beliefs are over stochastic adversaries, i.e., the agents believe that a given treatment yields the same outcome distribution in all agents. [^18] Agents have misspecified beliefs, in the sense that their support does not necessarily include the actual adversary. In particular, the actual adversary need not be stochastic. Misspecified beliefs are studied in a considerable recent literature in economic theory (starting from Berk,, 1966; Esponda and Pouzo,, 2016). However, this literature tends to focus on convergence and equilibria (e.g., Fudenberg et al.,, 2021; Esponda and Pouzo,, 2021; Esponda et al.,, 2021), whereas our paper takes the mechanism design perspective and is concerned with IPS-style statistical estimation. All agents share the same belief, a common simplification in economic theory. Such beliefs represent the current medical knowledge relevant to the clinical trial, or the learnings from user interactions collected by the online platform (which the platform may choose to share with its users).
Let us specify the agents’ beliefs formally. The outcome of each arm $a$ is an independent draw from some distribution $\psi^{*}(a)$ over outcomes each time this arm is chosen. For better notation, we identify states as mappings $\psi:[n]\to\Delta_{\Omega}$ from arms to distributions over outcomes, and $\psi^{*}$ as the true state. [^19] We use a standard notation: $\Delta_{\Omega}$ is the set of all distributions over $\Omega$. Then, the agents believe that $\psi^{*}$ is drawn from some Bayesian prior: distribution $\mathcal{P}$ over states. We posit that $\mathcal{P}$ is known to the mechanism, and assume that it has finite support.
@@
We ensure that each agent weakly prefers participation to the “outside option”. One way to formalize the outside option is that an agent could refuse to participate in the mechanism and instead choose an arm (e.g., a drug) that appears best based on the initial information. This achieves Bayesian-expected utility
$U_{\mathtt{out}}:=\max_{a\in[n]}\;\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\;\left[\,U_{\psi}(a)\,\right].$ (2.3)
More generally, we accommodates an arbitrary, exogenously given $U_{\mathtt{out}}\in[0,1]$. Such $U_{\mathtt{out}}$ could represent the Bayesian-expected utility of using a competing online platform.
@@
An arbitrary $U_{\mathtt{out}}\in[0,1]$ allows us to model several effects relevant to clinical RCTs. First, some medical treatments might not be available outside of a clinical trial. Then we could take the $\max$ in (2.3) only over the treatments that are. Conversely, a standard treatment might not be available inside the trial, e.g., in a new treatment vs. placebo trials. Then the $\max$ in (2.3) should also include the standard treatment. Finally, patients might assign positive utility to participation itself, e.g., because they value contributing to science or they expect to receive higher-quality and/or less expensive medical care. This can be modeled (albeit in a rather idealized way) by decreasing $U_{\mathtt{out}}$ in (2.3) by some known amount.